# Ground State Energy of Dilute Fermi Gases in 1D

### v1

@@ 1 Introduction
A well-known feature of ground state energies of interacting quantum gases is that they show a kind of universality in the interaction potential when the density of the gas is low. The best known example is perhaps the validity of the Lee–Huang–Yang energy expansion for 3D Bose gases with general potentials [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12], but other set-ups have been studied extensively in the mathematics literature in recent years, such as 2D dilute Bose gases [13, 14, 15, 16, 17, 18, 19] and 3D dilute Fermi gases [20, 21, 22, 23, 24, 25] (see the introductions of the papers cited for further references).
In contrast, this phenomenon had not been investigated much in one dimension. To address this, we previously studied the 1D Bose gas [26], and considered a gas of $N$ spinless bosons in the interval $[0,L]$ (with density $\rho=N/L$), interacting via the $N$-body Hamiltonian
@@
(where errors have been omitted to have an informal discussion; see [26, Theorem 1] for precise statements). This expression is in some sense universal because only the even wave scattering length appears to first order in the low-density ground state energy expansion, independent of the exact shape of the potential $v$. As indicated, a stronger universality (continuing the expansion to higher orders and finding that it depends only on $a$) has been demonstrated in 2 and 3D, but a crucial difference with higher dimensions is that in 1D, the leading-order term in the energy expansion is the ground state of the free Fermi gas. This means that the techniques used to prove (1.2) differ significantly from those used in 2 and 3D (see [26, Section 1.2] for a summary).
@@
The logical next question is: what happens for a dilute 1D spin-1/2 Fermi gas with Hamiltonian (1.1)? We can expect a free Fermi leading-order term like before, but this time, spin is present, and the wave function should be antisymmetric under the combined exchange of space and spin. If we again want to proceed by modifying the free Fermi state with 2-body scattering solutions, this means both even-wave and odd-wave scattering are needed: depending on the combined spin of the two particles involved in the scattering, the spatial wave function has to be symmetric or antisymmetric under exchange of the spatial coordinates. Indeed, in [26, Remark 5], we conjecture that the correct ground state energy is
@@ 2.1 The ground state energy of the dilute spin-$J$ Fermi gas
We consider a general even, compactly-supported, repulsive potential $v$ that is a measure. We the study the Hamiltonian
@@
We denote by $P_{S}$ is the projection onto the symmetric subspace $\mathbb{C}^{2J+1}\vee\mathbb{C}^{2J+1}$, while $P_{A}$ is the projection onto the antisymmetric subspace $\mathbb{C}^{2J+1}\wedge\mathbb{C}^{2J+1}$. We will add indices $i,j$ to indicate the spin projection is acting on particles $i$ and $j$, e.g. $P_{A}^{i,j}$. The even and odd-wave scattering lengths $a_{e}$ and $a_{o}$ were mentioned in (1.4) and (1.6) (see [26, Lemma 2 and Appendix A] for a precise definition, and Definition 8 below for a more general definition in the case of matrix-valued potentials).
@@
Let $v$ be a radial, repulsive interaction compactly supported in $R_{0},R_{0}]$, with $a_{e}$ and $a_{o}$ its even and odd wave scattering lengths. Consider a spin-$J$ Fermi gas with energy functional (2.2) and domain (2.3) restricted by any local boundary condition. Let $\rho=N/L$. Then, for $\rho|a_{e}|$, $\rho a_{o}$ and $\rho R_{0}$ sufficiently small, the ground state energy $E_{J}(N,L)$ satisfies
@@
where $\epsilon^{\textnormal{LS}}_{J}$ is the minimal energy per site of the spin-$J$ Lai–Sutherland model, i.e. the minimal expectation value of  [^1] We can assume periodic boundary conditions for the spin chain, but it does not matter for the theorem: other boundary conditions change $\epsilon^{\textnormal{LS}}_{J}$ by at most $N^{-1}$, which can be absorbed in the error.
@@
The assumption that the potential is compactly supported is not necessary: we may allow for potentials decaying fast enough to satisfy $\int_{b}^{\infty}v(x)x^{6+\epsilon}dx<\infty$ for some $b\geq 0$ and some $\epsilon>0$. In that case, the error in (2.4) is slightly more complicated. For simplicity, we only discuss the case of compactly-supported potentials in this paper, and refer to [26] for details on how to include potentials of non-compact support.
@@
Note that the notion of diluteness needed for the theorem is that $\rho|a_{e}|$, $\rho a_{o}$ and $\rho R_{0}$ are all small (where the absolute values are needed for $a_{e}$ because it can be negative, and the error involving $R_{0}$ cannot be avoided, see [26, Remarks 2 and 3]). Since the free case $v=0$ has $a_{e}=$-$\infty$ (see [26, Remark 4]), it cannot be considered dilute and the error $\rho|a_{e}|$ in the theorem is infinite.
@@
which proves a conjecture that we made in [26, Remark 5].
@@ 2.2 Ideas used in the proof
The most important ingredient in the proof are generalizations of the minimizations (1.4) for symmetric wave functions and (1.6) for antisymmetric wave functions. Those were used to prove the first-order ground state energy expansions for dilute spinless Bose and Fermi gases in [26]. For clarity, we first summarize the strategies used in the spinless case (see [26, Section 1.2] for a more extensive summary), followed by the adaptations needed for the case with spin.
The upper bound uses the trial state argument outlined in Section 1: for spinless bosons, the trial state departs from Girardeau’s solution of the impenetrable Bose gas (1.3), which is the free Fermi solution with absolute values to create a symmetric wave function. This produces the leading order in the expansion. The Girardeau wave function is then patched with the scattering solution (1.4) on some suitable small scale, and for the closest pair of particles only, which, up to negligible errors in the limits $\rho a_{e}\to 0$ and $\rho R_{0}\to 0$ (with $R_{0}$ the range of a compactly supported potential), gives our first-order result.
@@
The lower bound for spinless bosons in [26] again uses the minimization (1.4). The idea is that (1.4) can be written as
@@
for all $\psi\in H^{1}[-R,R]$ (note the boundary condition $\psi(R)=\psi(-R)=1$ in the minimization in (1.4)). This effectively lower bounds the kinetic and potential energy on [R,R][-R,R] by a symmetric delta potential at radius $R$. To generalize this to an $N$-boson wave function, we repeatedly apply (2.10) to obtain the symmetric delta potential for any neighbouring pairs of bosons. We then throw away the regions where $|x_{i+1}-x_{i}|\leq R$ (which gives a lower bound since $v$ is repulsive) and as a consequence, the two delta functions at radius $|x_{i+1}-x_{i}|=R$ collapse into a single delta at $|x_{i+1}-x_{i}|=0$, with value $4/(R-a_{e})$. This corresponds to the Lieb–Liniger energy of some wave function on the reduced interval, which is of course lower bounded by the explicitly known Lieb–Liniger ground state energy [28]. Heuristically, this corresponds to the following calculation: starting with an interval of length $L$, cut it back to length L(N1)RL-(N-1)R, so that the Lieb–Liniger ground state energy (see e.g. Lemma 27) with $c=2/(R-a_{e})$ and new density $N/(L-(N-1)R)=\rho(1+\rho R+\dots)$ gives a lower bound
@@
We also show that the ground state wave function has little weight in the regions that get thrown out, so that (2.11) is accurate.
@@ 2.4 Open problems
For example, in [26, Section 1.5], we conjectured that the energy expansions for dilute, spinless bosons and fermions (1.3) and (1.7) depend only on the scattering length to second order. In particular, for dilute bosons, we expect that the ground state energy is
@@
Of course, the other open problems for dilute, spinless 1D bosons mentioned in [26, Section 1.5]—studying momentum distributions and positive temperature for general potentials—extend to dilute gases with spin as well (see e.g. [61, 62, 63] in the physics literature). It would be particularly interesting to see to what extent the physics is still described by the Lai–Sutherland model (2.5).
@@ 3 Matrix-valued potentials
In the previous sections, we presented our results for scalar-valued potentials, but the upper bound we prove in the next section also applies to matrix-valued potentials. To state that result, we need a definition of the scattering length that is more general than the approach taken in (1.4) and (1.6) in the introduction (explained in more detail in [26, Lemma 2 and Appendix A], also see [64, Appendix C] for a 3D gas). Background facts needed for the following definition have been collected in Appendix A.
@@
where $a_{e}$ and $a_{o}$ are the even and odd-wave scattering lengths discussed in (1.4) and (1.6) in the introduction (see [26, Lemma 2 and Appendix A] for more properties).
@@ 4 Upper Bound for matrix-valued potentials
To state our upper bound in the most general way, we consider an even, repulsive potential $V$ that is a positive semidefinite matrix-valued measure acting on $\mathbb{C}^{2J+1}\otimes\mathbb{C}^{2J+1}$. We study the Hamiltonian
@@
We will decompose the potential as $V=vI+\tilde{V}$, where $v$ is a repulsive scalar potential as in the previous sections. For technical reasons, $\tilde{V}$ is assumed to be of bounded variation, for which we recall the following definition.
@@
where $\epsilon_{J}(\mathsf{A})$ is the ground state energy of the spin chain Hamiltonian  [^4] As indicated in Footnote 1, the boundary conditions of (4.5) are irrelevant for the statement. We will work with periodic boundary conditions for the spin chain in the proof.
@@
Let $v$ be a radial, repulsive interaction compactly supported in $R_{0},R_{0}]$, with $a_{e}$ and $a_{o}$ its even and odd wave scattering lengths. Consider a spin-$J$ Fermi gas with energy functional (2.2) and domain (2.3) with Dirichlet boundary conditions. Then, for $\rho R_{0}$, $\rho a_{e}$ and $\rho a_{o}$ sufficiently small, there exists a constant $C_{U}>0$ such that the ground state energy $E^{\textnormal{Dirichlet}}_{J}(N,L)$ satisfies
@@ 4.1 Preparatory facts about the free Fermi ground state
(See [26] for more details.) Wick’s theorem says that all free Fermi reduced densities matrices can be expressed in terms of $\gamma^{(1)}$. Also, we can verify that all derivatives of $\gamma^{(1)}$ can be bounded uniformly in the coordinates in terms of powers of $\rho$ (where the power can be obtained from dimensional analysis)—and this fact extends to all higher-order reduced density matrices by Wick’s theorem. The estimates above then follow from using the quotient rule for the fractions (if present) and Taylor expansion, simplified by symmetry and dimensional analysis to determine the correct power of $\rho$ (see [26, Lemma 10] for examples). ∎
@@ 4.3.2 Splitting the energy into parts and estimating them
The following lemmas show that the last four lines all end up in the error. The first three are proved in Section 4.3.4, the last (concerning the free Fermi gas) was proved in [26, Lemma 12].
@@ 4.3.3 A full trial state (Proof of Theorem 13)
as well as $\tilde{N}(\rho b)^{3}\leq\textnormal{const. }(\rho b)^{3/2}$, the desired result (4.4) follows (recall the remark about $\int\|\tilde{V}\|$ below (4.27)). Note the error $N\frac{\pi^{2}}{3}\rho^{2}\mathcal{O}(N^{-1})$ in (4.4) is only important when $N<(\rho b)^{-3/2}$, so it does not appear in (4.31). ∎

### v2

@@ 1 Introduction
A well-known feature of ground state energies of interacting quantum gases is that they show a kind of ‘universality’ in the interaction potential when the density of the gas is low. The best known example is perhaps the validity of the Lee–Huang–Yang energy expansion for 3D dilute Bose gases with general potentials [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12], but other set-ups have been studied extensively in the mathematics literature in recent years, such as 2D dilute Bose gases [13, 14, 15, 16, 17, 18] and 3D dilute Fermi gases [19, 20, 21, 22, 23, 24] (see the introductions of the papers cited for further references).
In contrast, in one dimension, dilute Bose and Fermi gases interacting with general potentials have not been investigated as much [^1] Note that weakly-interacting gases in 1D are not well described by the dilute limit, see Section 1.1 for a discussion.. To address this, we previously studied the 1D Bose gas [25], and considered a gas of $N$ spinless bosons in the interval $[0,L]$ (with density $\rho=N/L$), interacting via the $N$-body Hamiltonian
@@
(where error terms have been omitted to have an informal discussion; see [25, Theorem 1] for precise statements). This expression is in some sense ‘universal’ because only the even wave scattering length appears to first order in the low-density ground state energy expansion, independent of the exact shape of the potential $v$. As indicated, a stronger ‘universality’ (continuing the expansion to higher orders and finding that it depends only on $a$) has been demonstrated in 2 and 3D, but a crucial difference with higher dimensions is that in 1D, the leading-order term in the energy expansion is the ground state of the free Fermi gas. This means that the techniques used to prove (1.2) differ significantly from those used in 2 and 3D (see [25, Section 1.2] for a summary).
@@
The logical next question is: what happens for a dilute 1D spin-1/2 Fermi gas with Hamiltonian (1.1)? We can expect a free Fermi leading-order term like before, but this time, spin is present, and the wave function should be antisymmetric under the combined exchange of space and spin. If we again want to proceed by modifying the free Fermi state with 2-body scattering solutions, this means both even-wave and odd-wave scattering are needed: depending on the combined spin of the two particles involved in the scattering, the spatial wave function has to be symmetric or antisymmetric under exchange of the spatial coordinates. Indeed, in [25, Remark 5], we conjecture that the correct ground state energy is
@@
1.1 Luttinger theory and weakly-interacting 1D gases
This paper studies the ground state energy of dilute gases of particles with spin in 1D, and shows that the first correction term in the expansion only depends on the even and odd wave scattering lengths. This follows work that showed that for dilute gases in 2D and 3D, the ground state energy is even ‘universal’ for higher terms in the expansion (see the references at the start of the introduction).
Of course, the study of universal properties of 1D gases has a long history, in particular through Tomonaga–Luttinger theory, bosonization, and the Luttinger liquid universality class [53, 54, 55, 56, 57, 58, 59]. An important conclusion is that the theory of excitations of a large class of 1D models has a universal shape, independent of the exact microscopic model. However, this fact does not directly seem to comment on the ground state energy of the microscopic model, and how it depends on the shape of the interaction potential, which is the topic of interest here.
In the mathematical physics literature, many results have been obtained by Benfatto, Gallavotti, Mastropietro and others for weakly-interacting Fermi gases [60, 61, 62, 63, 64]. Using renormalization group techniques, these papers demonstrate a notion of universality, commonly used in the physics literature, that concerns the ground state energy, but also goes a great deal beyond that.
However, while in 2D and 3D the dilute and weakly-interacting limits are related, this is not the case in 1D. This is most easily seen in the Lieb–Liniger and Yang–Gaudin models, whose solution can entirely be expressed in terms of the parameter $\gamma=c/\rho$ (with $2c$ the parameter in front of the $\delta$-potential, see (1.12)) [27]. This means that the case of “fixed potential and low density” considered in this article corresponds to $\gamma\gg 1$, while the case of “fixed density and weak interactions” that is the focus of the references above corresponds to $\gamma\ll 1$. Hence, for the models with a $\delta$-interaction, the weakly-interacting regime and the low-density regime are opposites.
It is easy to carry out a similar scaling for potentials that are not concentrated at the origin like the $\delta$-function. In that case, the scaling that is needed to go between the two set-ups will also affect the range (or concentration) of the potential, but the conclusion remains the same: the regime considered in this article corresponds to strong interactions in the set-up used in the works by Benfatto, Gallavotti, Mastropietro and others, and so the results do not overlap.
@@ 2.1 The ground state energy of the dilute spin-$J$ Fermi gas
We consider a general even, compactly-supported, repulsive (i.e. positive) potential $v$ that is a measure. We the study the Hamiltonian
@@
We denote by $P_{S}$ is the projection onto the symmetric subspace $\mathbb{C}^{2J+1}\vee\mathbb{C}^{2J+1}$, while $P_{A}$ is the projection onto the antisymmetric subspace $\mathbb{C}^{2J+1}\wedge\mathbb{C}^{2J+1}$. We will add indices $i,j$ to indicate the spin projection is acting on particles $i$ and $j$, e.g. $P_{A}^{i,j}$. The even and odd-wave scattering lengths $a_{e}$ and $a_{o}$ were mentioned in (1.4) and (1.6) (see [25, Lemma 2 and Appendix A] for a precise definition, and Definition 8 below for a more general definition in the case of matrix-valued potentials).
@@
Let $v$ be an even, repulsive (i.e. positive) interaction compactly supported in $R_{0},R_{0}]$, with $a_{e}$ and $a_{o}$ its even and odd wave scattering lengths. Consider a spin-$J$ Fermi gas with energy functional (2.2) and domain (2.3) restricted by any Robin boundary condition. Let $\rho=N/L$. Then, for $\rho|a_{e}|$, $\rho a_{o}$ and $\rho R_{0}$ sufficiently small, the ground state energy $E_{J}(N,L)$ satisfies
@@
where $\epsilon^{\textnormal{LS}}_{J}$ is the minimal energy per site of the spin-$J$ Lai–Sutherland model, i.e. the minimal expectation value of  [^2] We can assume periodic boundary conditions for the spin chain, but it does not matter for the theorem: other boundary conditions change $\epsilon^{\textnormal{LS}}_{J}$ by at most $N^{-1}$, which can be absorbed in the error term.
@@
The assumption that the potential is compactly supported is not necessary: we may allow for potentials decaying fast enough to satisfy $\int_{b}^{\infty}v(x)x^{6+\epsilon}dx<\infty$ for some $b\geq 0$ and some $\epsilon>0$. In that case, the error term in (2.4) is slightly more complicated. For simplicity, we only discuss the case of compactly-supported potentials in this paper, and refer to [25] for details on how to include potentials of non-compact support.
@@
Note that the notion of diluteness needed for the theorem is that $\rho|a_{e}|$, $\rho a_{o}$ and $\rho R_{0}$ are all small (where the absolute values are needed for $a_{e}$ because it can be negative, and the error term involving $R_{0}$ cannot be avoided, see [25, Remarks 2 and 3]). Since the free case $v=0$ has $a_{e}=$-$\infty$ (see [25, Remark 4]), it cannot be considered dilute and $\rho|a_{e}|$ in the theorem is infinite.
@@
which proves a conjecture that we made in [25, Remark 5].
@@ 2.2 Ideas used in the proof
The most important ingredient in the proof are generalizations of the minimizations (1.4) for symmetric wave functions and (1.6) for antisymmetric wave functions. Those were used to prove the first-order ground state energy expansions for dilute spinless Bose and Fermi gases in [25]. For clarity, we first summarize the strategies used in the spinless case (see [25, Section 1.2] for a more extensive summary), followed by the adaptations needed for the case with spin.
The upper bound uses the trial state argument outlined in Section 1: for spinless bosons, the trial state departs from Girardeau’s solution of the impenetrable Bose gas (1.3), which is the free Fermi solution with absolute values to create a symmetric wave function. This produces the leading order in the expansion. The Girardeau wave function is then patched with the scattering solution (1.4) on some suitable small scale, and for the closest pair of particles only, which, up to negligible error terms in the limits $\rho a_{e}\to 0$ and $\rho R_{0}\to 0$ (with $R_{0}$ the range of a compactly supported potential), gives our first-order result.
@@
The lower bound for spinless bosons in [25] again uses the minimization (1.4). An important idea (see [2, 3] in the 3D case) is that (1.4) can be written as
@@
for all $\psi\in H^{1}[-R,R]$ (note the boundary condition $\psi(R)=\psi(-R)=1$ in the minimization in (1.4)). This effectively lower bounds the kinetic and potential energy on [R,R][-R,R] by a symmetric delta potential at radius $R$. To generalize this to an $N$-boson wave function, we repeatedly apply (2.10) to obtain the symmetric delta potential for any neighbouring pairs of bosons. Inspired by [70], we then throw away the regions where $|x_{i+1}-x_{i}|\leq R$ (which gives a lower bound since $v$ is positive) and as a consequence, the two delta functions at radius $|x_{i+1}-x_{i}|=R$ collapse into a single delta at $|x_{i+1}-x_{i}|=0$, with value $4/(R-a_{e})$. This corresponds to the Lieb–Liniger energy of some wave function on the reduced interval, which is of course lower bounded by the explicitly known Lieb–Liniger ground state energy [27]. Heuristically, this corresponds to the following calculation: starting with an interval of length $L$, cut it back to length L(N1)RL-(N-1)R, so that the Lieb–Liniger ground state energy (see e.g. Lemma 27) with $c=2/(R-a_{e})$ and new density $N/(L-(N-1)R)=\rho(1+\rho R+\dots)$ gives a lower bound
@@
As is done in [70], we also show that the ground state wave function has little weight in the regions that get thrown out, so that (2.11) is accurate.
@@ 2.4 Open problems
For example, in [25, Section 1.5], we conjectured that the energy expansions for dilute, spinless bosons and fermions (1.3) and (1.7) depend only on the scattering length to second order. In particular, for dilute bosons, we expect that the ground state energy is
@@
Of course, the other open problems for dilute, spinless 1D bosons mentioned in [25, Section 1.5]—studying momentum distributions and positive temperature for general potentials—extend to dilute gases with spin as well (see e.g. [73, 74, 75] in the physics literature). It would be particularly interesting to see to what extent the physics is still described by the Lai–Sutherland model (2.5).
@@ 3 Matrix-valued potentials
In the previous sections, we presented our results for scalar-valued potentials, but the upper bound we prove in the next section also applies to matrix-valued potentials. To state that result, we need a definition of the scattering length that is more general than the approach taken in (1.4) and (1.6) in the introduction (explained in more detail in [25, Lemma 2 and Appendix A], also see [76, Appendix C] for a 3D gas). Background facts needed for the following definition have been collected in Appendix A.
@@
where $a_{e}$ and $a_{o}$ are the even and odd-wave scattering lengths discussed in (1.4) and (1.6) in the introduction (see [25, Lemma 2 and Appendix A] for more properties).
@@ 4 Upper Bound for matrix-valued potentials
To state our upper bound in the most general way, we consider an even (repulsive) potential $V$ that is a positive semidefinite matrix-valued measure acting on $\mathbb{C}^{2J+1}\otimes\mathbb{C}^{2J+1}$. We study the Hamiltonian
@@
We will decompose the potential as $V=vI+\tilde{V}$, where $v$ is a positive scalar potential as in the previous sections. For technical reasons, $\tilde{V}$ is assumed to be of bounded variation, for which we recall the following definition.
@@
where $\epsilon_{J}(\mathsf{A})$ is the ground state energy of the spin chain Hamiltonian  [^5] As indicated in Footnote 2, the boundary conditions of (4.5) are irrelevant for the statement. We will work with periodic boundary conditions for the spin chain in the proof.
@@
Let $v$ be an even, positive interaction compactly supported in $R_{0},R_{0}]$, with $a_{e}$ and $a_{o}$ its even and odd wave scattering lengths. Consider a spin-$J$ Fermi gas with energy functional (2.2) and domain (2.3) with Dirichlet boundary conditions. Then, for $\rho R_{0}$, $\rho a_{e}$ and $\rho a_{o}$ sufficiently small, there exists a constant $C_{U}>0$ such that the ground state energy $E^{\textnormal{Dirichlet}}_{J}(N,L)$ satisfies
@@ 4.1 Preparatory facts about the free Fermi ground state
(See [25] for more details.) Wick’s theorem says that all free Fermi reduced densities matrices can be expressed in terms of $\gamma^{(1)}$. Also, we can verify that all derivatives of $\gamma^{(1)}$ can be bounded uniformly in the coordinates in terms of powers of $\rho$ (where the power can be obtained from dimensional analysis)—and this fact extends to all higher-order reduced density matrices by Wick’s theorem. The estimates above then follow from using the quotient rule for the fractions (if present) and Taylor expansion, simplified by symmetry and dimensional analysis to determine the correct power of $\rho$ (see [25, Lemma 10] for examples). ∎
@@ 4.3.2 Splitting the energy into parts and estimating them
The following lemmas show that the last four lines all end up in the error term. The first three are proved in Section 4.3.4, the last (concerning the free Fermi gas) was proved in [25, Lemma 12].
@@ 4.3.3 A full trial state (Proof of Theorem 13)
as well as $\tilde{N}(\rho b)^{3}\leq\textnormal{const. }(\rho b)^{3/2}$, the desired result (4.4) follows (recall the remark about $\int\|\tilde{V}\|$ below (4.27)). Note the error term $N\frac{\pi^{2}}{3}\rho^{2}\mathcal{O}(N^{-1})$ in (4.4) is only important when $N<(\rho b)^{-3/2}$, so it does not appear in (4.31). ∎